OpenStudy (ammarah):

Complete the square (find the value of n that makes each expression a a perfect square trinomial). 1. x^2 +14x+n

OpenStudy (ammarah):

@Luis_Rivera

OpenStudy (ammarah):

@.Sam.

OpenStudy (ammarah):

@wio

OpenStudy (bgrg007):

49

OpenStudy (raden):

that's right ^ (14/2)^2 = 7^2 = 49

OpenStudy (anonymous):

so then do u get it